OpenStudy (anonymous):

David had $4000 invested for 1/4 of a year at a simple interest rate of 6.1%. How much interest did he earn?

OpenStudy (abb0t):

I think you have confused what linear algebra is. This problem is algebra, not "linear algebra". To briefly explain, it is a branch of math that is concerned with vector spaces, often finite or countably infinite dimensions, as well as linear mappings between such spaces. You can learn more about it on Wikipedia, but it is a very tough course and requires some understanding of vector calculus. Now, moving onto your problem, you can use the formula: \(I=Prt\) to solve for \(I\). You are already given the inforamtion, P, r, and t. Plug it in and solve.

OpenStudy (abb0t):

P = what you invested r = rate t = time

OpenStudy (anonymous):

interest =prt/100 =(4000*1/4*6.1)/100 =61$